Description You are an internal auditor in a fund management company. You have been asked to verify the valuation of three bonds picked from a funds portfolio as of close of business on 31 December 2015 Bond A Face value Coupon Repayment Bullet at maturity Maturity Yield: $5,000,000 3.50% pa. payable semiannually 31 December 2020 The yield required by investors for a 5-year bond rated A+ by Standard & Poors and A1 by Moodys is 2.25% (semi-annual compounding) Bond B Face value Coupon Maturity Yield $4,000,000 Zero 31 December 2025 The yield required by investors for a 10-year bond rated AA- by Standard & Poors and Aa3 by Moodys is 5.25% (continuous compounding) Instructions Use a bond price calculator such as the rate function in Excel to calculate the value of the three bonds. Give brief details of how you calculated each answer What is the value of Bond A on 31 December 2015? What about the duration of Bond A? 2. What is the value of Bond B on 31 December 2015 What is the value of Bond A on 31 December 2015? What about the duration of Bond A? Responses What is the value of Bond B on 31 December 2015?
Answer #1

1. Value of Bond A : 87500* PVIFA(r,n) + 5000000*PVIF(r,n) = 170511 + 4584563 = 47,55,074.

2. Value of Bond B : Face value/ (1+r)n = 4000000/ (1+ 5.25%) 10 = 5,24,326.